You are considering a bond with a face value of $1 000 and a coupon rate of 2.0%. The bond has 16 year until maturity and coupon payments are paid semiannually. The yield to maturity on similar securities in the market is 8.3%

Given the information provided, what is the per period coupon payment for this bond?

What is the appropriate per period discount rate used to price this bond?

What is the current price of this bond?
Answer #1

Coupon per period =2%*1000/2 =10
Discount per period YTM =8.3%/2 =4.15%
Price of this bond =PV of Coupons+PV of Par value =10*(((1-(1+4.15%)^-32)/4.15%)+1000/(1+4.15%)^32 =447.58